# Smooth Tests for Normality in ANOVA

### v2

@@ 6.2 One-way random effects model
Now we turn to the one-way random effects model. The observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are generated from following equation:
@@
where the $\{A_{j}\}_{j=1}^{J}$ are i.i.d. random effects independent of the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, with $\mathrm{E}[A_{j}]=0$ and $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$. In this setting, we estimate $\sigma^{2}$ by
@@
where $\bar{Y}_{\cdot j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\varepsilon}_{ij}=Y_{ij}-\bar{Y}_{\cdot j}$. Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, leading to an identical expression for $\hat{\sigma}^{2}$ here as that in Section 3.2, too. Therefore, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are effectively eliminated by within-group centering via $Y_{ij}-\bar{Y}_{\cdot j}$.
6.3 Two-way models
Next, we extend the discussion to two-way ANOVA models. Suppose the observations
$\{Y_{ij_{1}j_{2}}\}_{i=1,j_{1}=1,j_{2}=1}^{N_{j_{1}j_{2}},J_{1},J_{2}}$ are generated from the two-way fixed effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.3)
@@ 6.3 Two-way models
Similarly, for the two-way random effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.4)
where the i.i.d. zero-mean random effects $\{A_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{B_{j_{2}}\}_{j_{2}=1}^{J_{2}}$, $\{D_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, are mutually independent. To test the normality of $e_{ij}$ in (6.4), one can mimic the strategy used in Section 6.2 to eliminate the random effects first through within-cell centering, and then construct the test statistic based on the resulting residuals.
To conclude, the methodology proposed in Section 3 can be seamlessly extended to more general ANOVA settings. This unified framework offers a flexible and effective approach to testing for normality in a wide range of linear models.
@@ 7 Concluding remarks
In this paper, we establish Neyman’s smooth tests for assessing the normality assumption in ANOVA models. For the proposed tests in three types of one-way fixed effects models, we demonstrate the asymptotic properties and validate the finite-sample performance through extensive simulations. Our approach provides a reliable tool for validating the normality assumption in diverse ANOVA settings.
Several directions remain open for future research. First, although we have developed tests within the ANOVA framework, the true data-generating mechanism may not always be known a priori. A promising approach is to conduct preliminary structural tests to identify the most suitable model before applying the proposed procedures. Second, the independence assumption on random errors can be relaxed to enable further theoretical analysis. For example, repeated clinical trials frequently exhibit autocorrelated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022). In such cases, our smooth test framework could potentially be adapted by incorporating estimated correlation structures. Finally, for the data-driven selection rule, it is intriguing to explore the case where the upper bound $D=D(N)\to\infty$ as $N\to\infty$ from the theoretical perspective.
@@ Appendix A Appendix
We provide the proof of our main theoretical results and some practical issues in this Appendix. The symbol $\lesssim$” means that the left side is bounded by a positive constant times the right side. The symbol $\asymp$” means that both sides are asymptotically equivalent.
@@
The proof is similar to that of Theorems 1, 2, and 3. Note that in the case of different means and same variance we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ without distinction of the null and alternatives. Our goal is to prove that
@@
Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
@@
Intuition behind the revised limiting null distribution+ (4.3).
We mainly follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case since the latter two cases can be derived analogously. Note that
$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)\\ =\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
The third term on the right-hand side of (A.30), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ). The event $\{\hat{K}=1\}$ is approximated by $\{N\hat{\Psi}^{2}_{1}-\log N\geq N\hat{\Psi}^{2}_{2}-2\log N\}=\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\hat{K}=2\}$ is approximated by $\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})>\log N\}$.
@@
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)\approx\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]\equiv H(x),\quad x\leq\log N,$
@@
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)\approx 2\left(1-\Phi\left(\sqrt{\log N}\right)\right)\approx 2\frac{\phi\left(\sqrt{\log N}\right)}{\sqrt{\log N}}=\sqrt{\frac{2}{\pi N\log N}},$
@@
Practical issues in simulation studies.
@@
Note that $c_{2k}=0$ when $k$ is odd and $c_{1k}=0$ when $k$ is even. Therefore, in practice, we only compute the non-zero elements based on the above two formulae for $k=1,\ldots,K$ with $K\leq 5$.

### v3

@@ 6.2 One-way random effects model
Next, we consider the one-way random effects model:
@@
where the random effects $\{A_{j}\}_{j=1}^{J}$ are i.i.d. with $\mathrm{E}[A_{j}]=0$, $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$, and are independent of the errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$. We estimate $\sigma^{2}$ by
@@
where $\bar{Y}_{\cdot j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\varepsilon}_{ij}=Y_{ij}-\bar{Y}_{\cdot j}$. Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, so the expression for $\hat{\sigma}^{2}$ is identical to that in Section 3.2. Consequently, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are removed by within-group centering.
6.3 Two-way ANOVA models
We now extend the discussion to two-way ANOVA models. Consider the two-way fixed effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.3)
@@ 6.3 Two-way ANOVA models
Similarly, consider the two-way random effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.4)
where the i.i.d. random effects $\{A_{j_{1}}\}$, $\{B_{j_{2}}\}$, $\{D_{j_{1}j_{2}}\}$ and the errors $\{e_{ij}\}$ are mutually independent. To test the normality of $e_{ij}$ in (6.4), one can apply the strategy of Section 6.2, first eliminating the random effects by within-cell centering and then constructing the test statistic based on the residuals.
To conclude, the methodology proposed in Section 3 can be seamlessly extended to more general ANOVA settings. This unified framework offers a flexible and effective approach to testing for normality across diverse linear models.
@@ 7 Concluding remarks
In this paper, we establish Neyman’s smooth tests for assessing the normality assumption in ANOVA models. For three types of one-way fixed effects models, we derive the asymptotic properties of the proposed tests and validate their finite-sample performance through extensive simulations. Our results provide a rigorous and practical tool for evaluating normality in a broad range of ANOVA settings.
Several directions remain open for future research. First, while the proposed tests are developed for specific ANOVA models, the true data-generating mechanism may not be known in practice. A natural extension is to combine our procedures with preliminary structural tests to identify the most appropriate ANOVA specification before applying the proposed methodology. Second, the assumption of independent random errors can be relaxed. For example, repeated clinical trials often involve correlated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022). In such cases, our smooth test framework may be adapted by incorporating estimated correlation structures. Finally, for the data-driven selection rule, it is of theoretical interest to explore the scenario where the upper bound $D=D(N)$ diverges with $N$, that is, $D\to\infty$ as $N\to\infty$.
@@ Appendix A Appendix
We provide the proof of our main theoretical results, along with some additional numerical results, in this Appendix. The symbol $\lesssim$” means that the left side is bounded by a positive constant times the right side. The symbol $\asymp$” means that both sides are asymptotically equivalent.
@@
The proof is similar to that of Theorems 1, 2, and 3. Note that in the case of different means and same variance, we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ without distinction of the null and alternatives. Our goal is to prove that
@@
Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in the fact that the first term of (A.4) should be further expressed as
@@
Intuition behind the revised limiting null distribution (4.3).
We primarily follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), and Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case, as the latter two cases can be derived analogously. Note that
$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)=$ $\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)$
$+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
The third term on the right-hand side of (A), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ). The event $\{\hat{K}=1\}$ is approximated by $\{N\hat{\Psi}^{2}_{1}-\log N\geq N\hat{\Psi}^{2}_{2}-2\log N\}=\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\hat{K}=2\}$ is approximated by $\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})>\log N\}$.
@@
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)$
$\approx$ $\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]\equiv H(x),\quad x\leq\log N,$
@@
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$\approx$ $2\left(1-\Phi\left(\sqrt{\log N}\right)\right)\approx 2\frac{\phi\left(\sqrt{\log N}\right)}{\sqrt{\log N}}=\sqrt{\frac{2}{\pi N\log N}},$
@@
Implement issues in simulation studies.
@@
Note that $c_{2k}=0$ when $k$ is odd and $c_{1k}=0$ when $k$ is even. Therefore, in our experiments, we only compute the non-zero elements based on the above two formulae for $k=1,\ldots,K$ with $K\leq 5$.
Additional simulation results.
The results of Experiment III are reported in Tables 14, 15, 16 and Figure 5. The main findings are broadly consistent with those in Experiment III, with two additional observations. First, Table 14 indicates that the limiting $\chi^{2}_{1}$ distribution for the data-driven test yields valid control of the Type I error. Second, the empirical distribution of $\hat{K}$ exhibits increased stability under both the null and the alternative hypotheses, with instances where it degenerates to a single value with probability one. Both phenomena are due to the increase in total sample size resulting from the larger number of groups, which is consistent with the theoretical results.
Table 14: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
10 0.028 0.046 0.062 0.036 0.054 0.060 0.048
20 0.046 0.066 0.078 0.052 0.068 0.056 0.052
30 0.050 0.056 0.044 0.040 0.048 0.044 0.042
40 0.054 0.054 0.042 0.040 0.044 0.058 0.052
50 0.052 0.062 0.052 0.056 0.054 0.048 0.042
60 0.062 0.050 0.060 0.048 0.048 0.052 0.046
70 0.068 0.052 0.052 0.054 0.044 0.058 0.054
80 0.052 0.034 0.044 0.062 0.052 0.048 0.046
90 0.048 0.048 0.034 0.050 0.058 0.042 0.040
100 0.064 0.048 0.052 0.048 0.052 0.058 0.054
110 0.044 0.048 0.058 0.058 0.032 0.060 0.056
120 0.052 0.058 0.066 0.044 0.052 0.080 0.078
130 0.046 0.050 0.038 0.058 0.062 0.058 0.052
140 0.058 0.046 0.032 0.054 0.046 0.056 0.050
150 0.046 0.040 0.060 0.050 0.054 0.038 0.038
Table 15: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.994 0.006 0 0 0
20 1 0 0 0 0
30 0.992 0.008 0 0 0
40 0.994 0.006 0 0 0
50 0.988 0.012 0 0 0
60 0.994 0.006 0 0 0
70 0.998 0.002 0 0 0
80 0.992 0.008 0 0 0
90 0.996 0.004 0 0 0
100 0.994 0.006 0 0 0
110 0.996 0.004 0 0 0
120 0.998 0.002 0 0 0
130 0.990 0.010 0 0 0
140 0.994 0.006 0 0 0
150 0.998 0.002 0 0 0
Table 16: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0 0 0 0.120 0.880
20 0 0 0 0.024 0.976
30 0 0 0 0 1
40 0 0 0 0 1
50 0 0 0 0 1
60 0 0 0 0 1
70 0 0 0 0 1
80 0 0 0 0 1
90 0 0 0 0 1
100 0 0 0 0 1
110 0 0 0 0 1
120 0 0 0 0 1
130 0 0 0 0 1
140 0 0 0 0 1
150 0 0 0 0 1
Figure 5: The sample means and error bars of $\hat{K}$ in Experiment III